### v4

# Similarity Field Theory: A Mathematical Framework for Intelligence

Abstract
We posit that persisting and transforming similarity relations form the structural basis of any comprehensible dynamic system. This paper introduces Similarity Field Theory, a mathematical framework that formalizes the principles governing similarity values among entities and their evolution. We define: (1) a similarity field $S:U\times U\to[0,1]$ over a universe of entities $U$, satisfying reflexivity $S(E,E)=1$ and treated as a directed relational field (asymmetry and non-transitivity are allowed); (2) the evolution of a system through a sequence $Z_{p}=(X_{p},\,S^{(p)})$ indexed by $p=0,1,2,\dots$; (3) concepts $K$ as entities that induce fibers $F_{\alpha}(K)=\{E\in U\mid S(E,K)\geq\alpha\}$, i.e., superlevel sets of the unary map $S_{K}(E):=S(E,K)$; and (4) a generative operator $G$ that produces new entities. Within this framework, we formalize a generative definition of intelligence: an operator $G$ is intelligent with respect to a concept $K$ if, given a system containing entities belonging to the fiber of $K$, it generates new entities that also belong to that fiber. Similarity Field Theory thus offers a foundational language for characterizing, comparing, and constructing intelligent systems. At a high level, this framework reframes intelligence and interpretability as geometric problems on similarity fields—preserving and composing level-set fibers—rather than purely statistical ones. We prove two theorems: (i) asymmetry blocks mutual inclusion; and (ii) stability implies either an anchor coordinate or asymptotic confinement to the target level (up to arbitrarily small tolerance). Together, these results constrain similarity-field evolution and motivate an interpretive lens that can be applied to large language models.
1 Introduction
We may regard the world as an evolving system. Within this system, similarity relations exist; without them, everything would collapse into undifferentiated chaos and the world we comprehend would cease to exist. Consider a cup before you: one second later it remains recognizably as a cup because certain similarities persist. If the cup falls and shatters, some similarities diminish while some others remain. These persisting and changing similarities jointly constitute the world we comprehend.
This paper formalizes this intuition by proposing a mathematical framework, Similarity Field Theory. Just as classical mechanics relies on calculus to analyze the continuous variation of physical quantities, Similarity Field Theory provides the essential constructs to analyze how intelligent systems preserve, modify, and generate similarity structures. We begin by establishing our foundational constructs—entities, concepts, and a similarity field. We then describe system evolution via a sequence $Z_{p}=(X_{p},S^{(p)})$ and introduce a generative operator, culminating in a formal, operational definition of intelligence as the capacity to preserve conceptual structure under generation.
Contributions.
To support rapid assessment by reviewers, we summarize core contributions.
  • A first-principles framework that treats graded, possibly asymmetric similarity as the primitive substrate for cognition and intelligence, with fibers as superlevel sets that define conceptual membership.
  • A sequence-based evolution formalism, including a strengthened stability theorem that characterizes anchors versus eventual confinement and an incompatibility theorem showing how asymmetry forbids reciprocal inclusion under cross-referenced thresholds.
  • An operational, generative definition of intelligence in terms of fiber preservation and coverage/fidelity measures.
  • A framework for interpretability that views networks as compositions of calibrated similarity fields.
2 Related Work
Similarity Field Theory builds upon and extends several foundational areas. We position it as a first-principles framework that elevates similarity to a primitive substrate, complementing and unifying existing concepts rather than deriving from any single field.
2.1 Fuzzy Set Theory: A Formal Analogy and an Essential Distinction
A core structure in our framework—a grouping of entities based on their similarity to a given concept—is formally analogous to the $\alpha$-cut’ in Zadeh’s Fuzzy Set Theory [18]. This formal link provides a useful bridge to a mature field. However, this similarity is formal rather than foundational. The primary goal of Fuzzy Set Theory is to represent static knowledge under conditions of vagueness. In contrast, the central aim of Similarity Field Theory is to formulate a theory of intelligence based on the dynamics of how conceptual structures evolve and are generated, which can reuse fuzzy constructs when convenient, without being reducible to them. Our fundamental primitive is the similarity relation itself, modeled as a directed field $S$. Our framework is therefore not an extension of Fuzzy Set Theory but a relation-centric dynamic similarity theory.
2.2 Metric Spaces: Relaxed Axioms for Cognitive Modeling
Our framework deliberately departs from the rigid axioms of metric spaces. The requirement for a distance function to be symmetric and to obey the triangle inequality, while essential for describing physical space, is often too restrictive for modeling the nuances of cognitive or conceptual space. In practice, machine learning routinely employs measures such as cosine similarity, KL divergence, Jensen–Shannon divergence, and general Bregman divergences; several of these are asymmetric and/or violate the triangle inequality, motivating a relaxation of metric axioms for conceptual geometry [15, 9, 11, 4, 2, 1, 17]. For the following examples, we fix a single similarity notion $S$ over the shared universe $U$ (e.g., cosine in a common embedding space or a corpus-based associative measure); all pairwise comparisons refer to this same $S$. Consider lion (A), domestic cat (B), and household pet (C): $S(A,B)$ is high (both are felids), $S(B,C)$ is high (cats are prototypical pets), yet $S(A,C)$ is much lower (lions are not household pets), yielding high $S(A,B)$ and $S(B,C)$ but comparatively low $S(A,C)$.
2.3 Category Theory: A Language of Structure
Category theory provides a language for abstract structures, describing objects and their connections (morphisms). In the standard view, the link between two objects is a discrete, binary matter: a morphism either exists or does not. While powerful, this is insufficient to capture the graded nature of conceptual relations. We use a similarity field $S$ taking values in $[0,1]$ to quantify the strength of belonging/connection; given a concept $K$, we express structure and invariants via threshold sets of $S(\cdot,K)$; generation is “intelligent” iff it preserves and expands these threshold sets along the evolution. We do not read $S$ as a discrete morphism between arbitrary objects; instead, we center on the concept $K$ and study $S(\cdot,K)$ and its fibers $F_{\alpha}(K)$, using strength rather than mere existence to characterize belonging and connection. More importantly, this framework targets any system organized by similarity—from perception and learning to machine models, social bargaining, and institutional design—rather than being confined to categorical settings or specialist uses of “fiber”. Enriched categorical phrasing is only an explanatory lens; our theory and results hold even without adopting it [10, 8].
2.4 A Theoretical Foundation for Model Interpretability
2.4.1 Current Explorations and the Need for a Theory
The academic community has made explorations into model interpretability, developing techniques to trace information “circuits” [12], identify “persona vectors” in activation space [3, 6], and perform “representation engineering” [19, 16, 14]. These empirical works enhance our ability to inspect AI systems, but they often lack a unifying mathematical theory that explains how these observed structures emerge from first principles. The Similarity Field Theory framework is proposed not to replace these techniques, but to offer a foundational, geometric mathematical perspective: it casts models as calibrated similarity fields whose level-set fibers constitute concepts, thereby unifying and guiding interpretability studies at a deeper structural level beyond purely statistical views.
2.4.2 Beyond a Purely Statistical View
Viewing a neural network as a statistical learning machine is a dominant and effective practice; it excels at characterizing in-distribution input–output behavior but offers limited insight into the formation of internal conceptual structure and into out-of-distribution (OOD) phenomena. Similarity Field Theory complements this view from the angle of conceptual similarity: we regard a trained model as a complex system composed of nested similarity fields and restate interpretability as a task of conceptual deconstruction—identifying latent concepts formed inside the model and defining them via their fibers (the sets of inputs that elicit high similarity responses). Consequently, the research goal shifts from merely fitting external statistics to understanding and organizing these concept fibers, enabling structured judgments beyond the training distribution in accordance with conceptual similarity rather than being constrained by frequency statistics of the training distribution. For example, if the training data include only “red circles” and “green triangles” and never “green circles,” the model can still classify a green circle as a circle because it matches along the fiber of the concept “circle,” rather than relying on frequency co-occurrence.
Box 1  |  Why fiber-based structure can generalize when joint-frequency MLE collapses Consider two attributes—color $C\in\{\mathrm{red},\mathrm{green}\}$ and shape $H\in\{\mathrm{circle},\mathrm{triangle}\}$—and entities of the form $(C,H)$. Suppose the training set contains only $(\mathrm{red},\mathrm{circle})$ and $(\mathrm{green},\mathrm{triangle})$. A joint-frequency maximum-likelihood estimate assigns $\widehat{P}_{\mathrm{MLE}}(C=\mathrm{green},H=\mathrm{circle})=\frac{n(\mathrm{green},\mathrm{circle})}{N}=0,$ so any decision rule that relies on the joint MLE as “support” cannot endorse the unseen combination $(\mathrm{green},\mathrm{circle})$ without introducing extra structure. In Similarity Field Theory (SFT), concepts are modeled as superlevel sets of a similarity field. Let $K$ denote the concept “circle”, and define the fiber $F_{\alpha}(K)=\{E:\ S(E,K)\geq\alpha\}$. Let $X_{K}$ be a small exemplar set intended to lie in the fiber; in the simplest case, $X_{K}=\{x\}$ with $x=(\mathrm{red},\mathrm{circle})$. Define an exemplar-to-concept readout $\hat{S}(E,K)=\max_{x\in X_{K}}\min\big\{S_{K}(E,x),\,S(x,K)\big\},$ where $S_{K}(\cdot,\cdot)$ denotes a $K$-aligned similarity probe (e.g., a shape-focused similarity). Lemma (one-step fiber extension). If there exists $x\in X_{K}$ such that $S(x,K)\geq\alpha$ and $S_{K}(E,x)\geq\alpha$, then $\hat{S}(E,K)\geq\alpha$ and hence $E\in\hat{F}_{\alpha}(K)$. Thus, even when $n(\mathrm{green},\mathrm{circle})=0$, the unseen entity $E=(\mathrm{green},\mathrm{circle})$ can be admitted into the “circle” fiber whenever it is sufficiently similar (in the $K$-aligned sense) to an exemplar already inside the fiber. By contrast, additive smoothing (e.g., Laplace/Dirichlet) redistributes mass to unseen cells without preferentially selecting $(\mathrm{green},\mathrm{circle})$ over other unobserved combinations unless additional factorization or concept-specific structure is imposed.
3 Foundational Constructs
Definition 1 (Entity).
An entity, denoted $E$, is any identifiable element of discourse. An entity can be atomic, a concept, or itself a similarity relation.
Definition 2 (Universe).
The universe, denoted $U$, is the set of all entities.
Definition 3 (Concept).
A concept, denoted $K$, is an entity that serves as a representation for a class of other entities. It induces a unary map $S_{K}(E):=S(E,K)$ and thereby defines fibers as its superlevel sets.
Definition 4 (Similarity Field).
A similarity field is a mapping $S:U\times U\to[0,1]$ satisfying reflexivity: $S(E,E)=1$ for all $E\in U$. We treat $S$ as a directed relational field; symmetry and transitivity are not assumed.
Definition 5 (Fiber).
Given a concept $K\in U$ and a threshold $\alpha\in[0,1]$, the fiber induced by $K$ is the superlevel set:
$F_{\alpha}(K)=\{E\in U\mid S(E,K)\geq\alpha\}.$
Structural assumptions (minimal). We work with the weakest topological or graph structures needed: a domain of observables $D\subseteq[0,1]^{n}$ admits well-defined limits and closures (e.g., by restricting to compact subsets), so that sequence limits, $\omega$-limit sets, and level sets are well-defined. Smoothness is assumed only when explicitly required.
Terminology clarification. Throughout, “field” means a value-assigning function in the geometric sense; it is neither an algebraic field nor a physical field theory.
3.1 On the Nature of the Similarity Field
It is a crucial feature of our framework that the similarity field $S$ is defined broadly, requiring only reflexivity. We deliberately do not enforce two stronger conditions common in classical mathematics:
  • Symmetry: We do not assume that $S(E_{1},E_{2})=S(E_{2},E_{1})$. This allows the framework to model asymmetric cognitive phenomena in reality. Classic work in cognitive psychology has shown that people often judge, for example, “North Korea is similar to China” as more natural, and endorse it more frequently, than the reversed statement “China is similar to North Korea”, even though the underlying pair of countries is the same; such reference-point and salience effects systematically break symmetry in human similarity judgments [17].
  • Transitivity: High similarity need not be transitive. In conceptual space, consider lion (A), domestic cat (B), and household pet (C): $S(A,B)$ is high (both are felids), $S(B,C)$ is high (domestic cats are prototypical pets), yet $S(A,C)$ is much lower (lions are not household pets), yielding high $S(A,B)$ and $S(B,C)$ but comparatively low $S(A,C)$. This pattern departs from metric triangle-inequality assumptions and matches empirical observations about graded category structure and conceptual similarity [17, 5].
This deliberate relaxation is motivated by the arguments presented in the main text, distinguishing Similarity Field Theory from frameworks based on Metric Spaces or strict Equivalence Relations. While a field satisfying all three properties is a valid special case, our broader definition is essential for modeling the continuous and overlapping nature of conceptual structures in the real world.
3.2 Properties of Similarity Fields
The properties of the similarity field, particularly its allowance for asymmetry, give rise to non-trivial principles governing conceptual membership. We formalize some of these properties below.
Proposition 1 (Fiber monotonicity).
Fix $K\in U$. If $\alpha\geq\beta$, then $F_{\alpha}(K)\subseteq F_{\beta}(K)$.
Proof.
If $E\in F_{\alpha}(K)$, then $S(E,K)\geq\alpha\geq\beta$, hence $E\in F_{\beta}(K)$. ∎
Proposition 2 (Intersection identity).
For fixed $K\in U$ and thresholds $\{\alpha_{r}\}_{r=1}^{m}$, we have $\bigcap_{r=1}^{m}F_{\alpha_{r}}(K)=F_{\max_{r}\alpha_{r}}(K)$.
Proof.
If $E\in\bigcap_{r=1}^{m}F_{\alpha_{r}}(K)$, then $S(E,K)\geq\alpha_{r}$ for all $r$, so $S(E,K)\geq\max_{r}\alpha_{r}$ and $E\in F_{\max_{r}\alpha_{r}}(K)$. The reverse inclusion is immediate. ∎
Proposition 3.
If $S$ is a similarity field, d=1Sd=1-S is not.
Proof.
For $d$ to be a similarity field, it must satisfy $d(E,E)=1$. By definition, d(E,E)=1S(E,E)=11=0d(E,E)=1-S(E,E)=1-1=0. This implies the contradiction $0=1$. ∎
Proposition 4 (Closure under Multiplication).
If $S_{1},S_{2}$ are similarity fields, so is their pointwise product $S_{1}\cdot S_{2}$.
Proof.
The range remains $[0,1]$. For reflexivity, $(S_{1}\cdot S_{2})(E,E)=S_{1}(E,E)\cdot S_{2}(E,E)=1\cdot 1=1$. ∎
Proposition 5 (Closure under Convex Combination).
If $\{S_{k}\}$ is a collection of similarity fields and $\{w_{k}\}$ are non-negative weights summing to 1, then $\sum w_{k}S_{k}$ is a similarity field.
Proof.
The range remains $[0,1]$. For reflexivity, $(\sum w_{k}S_{k})(E,E)=\sum w_{k}S_{k}(E,E)=\sum w_{k}\cdot 1=1$. ∎
Theorem 1 (Incompatibility Theorem).
Given two distinct entities $E_{1},E_{2}\in U$. Let their similarity values be $x=S(E_{1},E_{2})$ and $y=S(E_{2},E_{1})$, produced by the single similarity field $S$. If $x\neq y$ (i.e., the similarity relation is asymmetric), then the following two conditions cannot both be true simultaneously:
  1. 1.
    $E_{1}$ belongs to the fiber induced by $E_{2}$ as a concept, with threshold $y$ (i.e., $E_{1}\in F_{y}(E_{2})$).
  2. 2.
    $E_{2}$ belongs to the fiber induced by $E_{1}$ as a concept, with threshold $x$ (i.e., $E_{2}\in F_{x}(E_{1})$).
Proof.
We use proof by contradiction. Assume that both condition 1 and condition 2 are simultaneously true.
  • From condition 1 being true, by the definition of a fiber $F_{y}(E_{2})=\{E\in U\mid S(E,E_{2})\geq y\}$, we must have $S(E_{1},E_{2})\geq y$. Substituting $x=S(E_{1},E_{2})$, we get $x\geq y$.
  • From condition 2 being true, by the definition of a fiber $F_{x}(E_{1})=\{E\in U\mid S(E,E_{1})\geq x\}$, we must have $S(E_{2},E_{1})\geq x$. Substituting $y=S(E_{2},E_{1})$, we get $y\geq x$.
Combining these two results, we have $x\geq y$ and $y\geq x$, which necessarily implies $x=y$. However, this contradicts the theorem’s premise that $x\neq y$. Therefore, our initial assumption is false, meaning that condition 1 and condition 2 cannot both be true simultaneously. ∎
Significance of Theorem 1: The key point is that asymmetry induces a one-sided constraint: under cross-referenced thresholds, reciprocal inclusion becomes impossible unless the similarity relation is symmetric. This highlights a structural constraint: asymmetry can impose inherent limitations and a forced imbalance on conceptual membership. It establishes the impossibility of reciprocal inclusion: if $E_{1}$’s perspective on $E_{2}$ differs from $E_{2}$’s perspective on $E_{1}$ (i.e., $x\neq y$), then they cannot simultaneously be members of each other’s conceptual fibers when the threshold for inclusion is set by the other’s standard. In a stylized negotiation setting where each party sets its acceptance threshold using the other party’s valuation, the theorem yields a simple infeasibility condition: if the assessments are asymmetric, mutual agreement cannot be achieved under these cross-referenced thresholds unless at least one party adjusts its acceptance criterion, for example by slightly lowering a price (or raising an offer), which effectively shifts the fiber threshold and may, when the parties are near the boundary, turn non-membership into membership.
This principle is illustrated by the Novice ($N$) and Expert ($E$) scenario. Suppose $S(N,E)=0.9$ and $S(E,N)=0.2$. The Novice may belong to the Expert’s fiber ($N\in F_{0.2}(E)$ since $0.9\geq 0.2$), but the theorem guarantees the impossibility of the reverse: the Expert cannot belong to the Novice’s fiber at the Novice’s standard ($E\notin F_{0.9}(N)$ since $0.2<0.9$). The asymmetry in similarity mathematically forbids a symmetric sense of belonging.
Party AParty B "My position is worth 0.3 to them." $S(B,A)=0.3$ "My position is worth 0.8 to them." $S(A,B)=0.8$ Party A’s value (0.8) is high enough. Agreement PossibleParty B’s value (0.3) is too low. Agreement Fails
Figure 1: The Incompatibility Theorem visualized through a negotiation. Each party’s "acceptance fiber" is defined by the other’s perceived value. Party B’s offer ($S(B,A)=0.3$) is too low to enter Party A’s acceptance fiber (threshold 0.8). Conversely, Party A’s offer ($S(A,B)=0.8$) easily clears Party B’s threshold (0.3). Because the condition is not met for both, mutual agreement is impossible.
4 System Evolution via Sequences
Sequence-only stance. We do not introduce flow maps; all notions of evolution, stability, and $\omega$-limits are expressed purely in terms of sequences.
Definition 6 (System-State Sequence).
A system-state sequence is a sequence $Z_{p}=(X_{p},S^{(p)})$ with $p=0,1,2,\dots$, where $X_{p}\subseteq U$ is a finite set of entities and $S^{(p)}:U\times U\to[0,1]$ is a similarity field at step $p$.
Index interpretation. The index $p$ may denote discrete time, training steps/epochs, rounds of interaction, or any other well-ordered design parameter; our results depend only on its ordering, not on its physical interpretation. Even if an underlying continuous time exists, empirical measurements are effectively discrete at finite resolution; accordingly, we adopt a sequence-based formalism while remaining compatible with continuous-time models where available.
The evolution of a system-state sequence can be characterized by two primary modes:
  1. 1.
    Field Variation: For a fixed set of entities, their relational structure changes along the sequence via $S^{(p)}(E_{i},E_{j})$. When $p$ samples a continuous underlying parameter, differentiability of $S^{(p)}$ may be assumed only when needed.
  2. 2.
    Discrete Composition Change: The set $X_{p}$ evolves via the addition/removal/modification of entities. This mode is driven by a generative operator (defined later).
Eventual confinement and absorbing sets. A subset $\mathcal{A}\subseteq U$ is called an absorbing set for a given evolution sequence if there exists $p^{\star}$ such that for all $p\geq p^{\star}$, the entities under consideration lie in $\mathcal{A}$ and never leave it along the subsequent sequence. When $\mathcal{A}$ is a fiber $F_{\alpha}(K)$, we may say the sequence exhibits eventual confinement inside the fiber of concept $K$.
4.1 The Principle of Foundational Stability
Background and Core Significance.
This theorem establishes, with mathematical rigor, a foundational principle of cognition that scales from the individual mind to the macro-society: any coherent, stable cognitive system is mathematically dependent on at least one long-term, foundational belief or concept that serves as a source of stability. Without a stability source, collapse is inevitable. In any system composed of interconnected elements, the stability of one property necessarily depends on the stability of its determinants.
Topological convention.
We fix an embedding $D\subseteq[0,1]^{n}$ and endow $D$ with its subspace topology. Unless otherwise stated, closures $\overline{A}$ appearing in $C_{c}^{\star}$ are taken in the ambient space $0,1]^{n}$. The $\omega$-limit set collects cluster points in the ambient closure $\overline{D}$ [7].
Set-up.
Let $f:D\to[0,1]$ be non-constant. For each index $p$, define the determining vector
$\mathbf{v}_{p}=\bigl(S^{(p)}(E_{1},E_{1}^{\prime}),\ldots,S^{(p)}(E_{n},E_{n}^{\prime})\bigr)\in D,\qquad y_{p}:=f(\mathbf{v}_{p}).$
A sequence $y_{p}$ is called stable if there exists $c\in[0,1]$ such that for every $\epsilon>0$ there is $p^{\ast}$ with $|y_{p}-c|\leq\epsilon$ for all $p\geq p^{\ast}$. For $c\in[0,1]$ and $\epsilon>0$, define the level-set tube $N_{\epsilon}(c):=\{x\in D:\ |f(x)-c|\leq\epsilon\}$ and the closed level set
$C_{c}^{\star}:=\bigcap_{m=1}^{\infty}\overline{\{x\in D:\ |f(x)-c|\leq 1/m\}}.$
The $\omega$-limit set is $\omega(\mathbf{v}):=\{x\in\overline{D}:\ \exists\ p_{k}\to\infty\ \text{with}\ \mathbf{v}_{p_{k}}\to x\}$. A coordinate sequence $\pi_{i}(\mathbf{v}_{p})$ is called an anchor if it converges.
Theorem 2 (Stability Theorem, strengthened).
Assume $y_{p}=f(\mathbf{v}_{p})$ is stable with limit $c$. Then:
  1. 1.
    For every $\epsilon>0$ there exists $p_{\epsilon}$ such that $\mathbf{v}_{p}\in N_{\epsilon}(c)$ for all $p\geq p_{\epsilon}$. Hence $\omega(\mathbf{v})\subseteq C_{c}^{\star}$.
  2. 2.
    If $f$ is continuous, then $\omega(\mathbf{v})\subseteq f^{-1}(\{c\})$.
  3. 3.
    Either some coordinate $\pi_{i}(\mathbf{v}_{p})$ converges, or $\omega(\mathbf{v})$ contains at least two distinct points and is contained in $C_{c}^{\star}$.
Lemma 1 (Separation Criterion).
Let $A,B\subseteq D$ and $\delta>0$ be such that $\mathbf{v}_{p}$ visits $A$ and $B$ infinitely often, and $|f(a)-f(b)|\geq\delta$ for all $a\in A$, $b\in B$. Then $y_{p}=f(\mathbf{v}_{p})$ does not converge.
Proof.
Choose subsequences along visits to $A$ and to $B$. Their images under $f$ differ by at least $\delta$, so $\limsup y_{p}-\liminf y_{p}\geq\delta>0$, which rules out convergence. ∎
Lemma 2.
If $\{\mathbf{v}_{p}\}$ has a unique cluster point $x^{\star}$ in the ambient compact space $0,1]^{n}$ (hence also in $\overline{D}$), then $\mathbf{v}_{p}\to x^{\star}$, so each coordinate $\pi_{i}(\mathbf{v}_{p})$ converges. Consequently, if no coordinate converges, then the set of ambient cluster points contains at least two distinct points.
Proof.
If $\mathbf{v}_{p}\nrightarrow x^{\star}$, there exist $\eta>0$ and infinitely many $p$ with $\|\mathbf{v}_{p}-x^{\star}\|\geq\eta$. By compactness of $0,1]^{n}$, this subsequence has a convergent sub-subsequence with limit $y\neq x^{\star}$, contradicting uniqueness. ∎
Proof of Theorem 2.
Let $y_{p}=f(\mathbf{v}_{p})$ and assume $y_{p}\to c$. Then for every $\epsilon>0$ there exists $p_{\epsilon}$ such that $|f(\mathbf{v}_{p})-c|\leq\epsilon$ for all $p\geq p_{\epsilon}$, i.e., $\mathbf{v}_{p}\in N_{\epsilon}(c)$ eventually. Hence any cluster point $x$ of $\{\mathbf{v}_{p}\}$ lies in every $\overline{N_{1/m}(c)}$, and thus $\omega(\mathbf{v})\subseteq C_{c}^{\star}$. If, in addition, $f$ is continuous and $\mathbf{v}_{p_{k}}\to x\in\omega(\mathbf{v})$, then $f(x)=\lim_{k\to\infty}y_{p_{k}}=c$, so $\omega(\mathbf{v})\subseteq f^{-1}(\{c\})$. For item (3), if some coordinate $\pi_{i}(\mathbf{v}_{p})$ converges, we are in the first alternative of item (3). Otherwise, by Lemma 2, the ambient cluster set contains at least two distinct points, and by the first paragraph $\omega(\mathbf{v})\subseteq C_{c}^{\star}$, which is the second alternative of item (3). ∎
Equivalent tail characterization.
For sequences $y_{p}=f(\mathbf{v}_{p})$, $y_{p}\to c$ if and only if for every $\epsilon>0$ there exists $p_{\epsilon}$ such that $\mathbf{v}_{p}\in N_{\epsilon}(c)$ for all $p\geq p_{\epsilon}$.
Proof.
If $y_{p}\to c$, then for every $\epsilon>0$ there exists $p_{\epsilon}$ with $|f(\mathbf{v}_{p})-c|\leq\epsilon$ for all $p\geq p_{\epsilon}$, hence $\mathbf{v}_{p}\in N_{\epsilon}(c)$. Conversely, if for every $\epsilon>0$ there exists $p_{\epsilon}$ with $\mathbf{v}_{p}\in N_{\epsilon}(c)$ for all $p\geq p_{\epsilon}$, then $|f(\mathbf{v}_{p})-c|\leq\epsilon$ for all large $p$, so $y_{p}\to c$. ∎
Corollary 1 (Contrapositive Test).
If there exists $\epsilon_{0}>0$ such that $\mathbf{v}_{p}\notin N_{\epsilon_{0}}(c)$ for infinitely many $p$, then $y_{p}$ cannot be stable at $c$.
Sufficient tail condition.
For every $\epsilon>0$, if there exists $p_{\epsilon}$ with $\mathbf{v}_{p}\in N_{\epsilon}(c)$ for all $p\geq p_{\epsilon}$, then $y_{p}\to c$, and in particular $\omega(\mathbf{v})\subseteq C_{c}^{\star}$. In this sense, an absorbing set is any set that the sequence eventually enters and then remains in.
Proof.
The assumption implies $|f(\mathbf{v}_{p})-c|\leq\epsilon$ for all large $p$ and every $\epsilon>0$, hence $y_{p}\to c$. Any cluster point must lie in every $\overline{N_{1/m}(c)}$, so $\omega(\mathbf{v})\subseteq C_{c}^{\star}$. ∎
Illustrative Examples.
Mutual cancellation: for $f(v_{1},v_{2})=(v_{1}+v_{2})/2$ with $v_{1}(p)=0.5+0.5\sin p$ and $v_{2}(p)=0.5-0.5\sin p$, each coordinate oscillates but $f\equiv 0.5$ is perfectly stable. Saturation: for $f(x)=\sigma(w^{\top}x+b)$ with a bounded monotone activation $\sigma$, if $w^{\top}x+b$ remains in a saturated regime, then $f(x)$ is nearly constant despite input jitter.
5 Generative Dynamics and Intelligence
Definition 7 (Generative Operator).
A generative operator, denoted $G$, is a function that maps a current set of entities to a new set:
$E_{\text{new}}=G(X_{p}).$
The system then undergoes discrete evolution according to the update rule $X_{p+1}=X_{p}\cup E_{\text{new}}$, yielding an updated pair $Z_{p+1}=(X_{p+1},S^{(p+1)})$.
In physics, energy is the ability to do work. Analogously, our prior work [13] gave an informal definition of intelligence: “intelligence is the ability, given entities exemplifying a concept, to generate entities exemplifying the same concept.” This conceptual work is translated into the physical world by the action of a generative operator. For example, a large language model that, when given the conceptual requirement “a Python function to efficiently sort a list of numbers”, generates the correct and functional source code.
Within the Similarity Field Theory framework, this principle is expressed formally. The premise, “given entities exemplifying a concept,” corresponds to a system-state $Z_{p}$ that contains a subset of entities $X_{K}\subseteq X_{p}$, where every entity in $X_{K}$ belongs to a fiber $F_{\alpha}(K)$.
Intelligence with respect to a concept $K$ is the capacity of a generative operator $G$ to produce new entities $E^{\prime}\in G(X_{p})$ that also belong to the fiber $F_{\alpha}(K)$ by satisfying the condition:
$S(E^{\prime},K)\geq\alpha.$
Example (ChatGPT, Ghibli style). Let $K$ denote the concept “Studio Ghibli–style image.” Provide a reference context of Ghibli-style images and prompt ChatGPT to produce new frames under the same context. Using a fixed similarity function $S(\cdot,K)$ (e.g., style embeddings or calibrated ratings) to score each output, if a substantial portion of the outputs satisfy $S\geq\alpha$ (high coverage) and the average score is high (high fidelity), then ChatGPT is $\alpha$-intelligent with respect to the concept “Ghibli-style image.”
We can therefore quantify intelligence by within-output coverage (fraction in $F_{\alpha}(K)$) and fidelity (mean $S(E^{\prime},K)$).
5.1 Learning
Learning is the process by which a system’s generative operator $G$ or its similarity field $S^{(p)}$ is modified over the sequence (i.e., as $p$ increases) to improve its intelligence with respect to a concept $K$. A learning rate can be conceptualized as the rate of change in the fidelity of generated entities, such as the difference quotient for $\mathbb{E}[\,S^{(p)}(G(X_{p}),K)\,]$ along $p$.
5.2 Creativity as Re-Contextualization
Furthermore, our framework provides a formal perspective on creativity. This is not treated as creation ex nihilo, but as the formation of a new concept initiated by the re-contextualization of an existing entity. This occurs when an entity $E$, already a member of a source concept $K_{\text{source}}$, is discovered to also possess high similarity to a different, nascent target concept $K_{\text{target}}$. Prior to this discovery, the fiber of the target concept, $F_{\beta}(K_{\text{target}})$, might have been empty within the system. The act of recognizing that $S(E,K_{\text{target}})\geq\beta$ populates this fiber for the first time, establishing $K_{\text{target}}$ as a valid concept and positioning $E$ as its foundational instance. The history of WD-40 illustrates this. Its formula ($E$) was an entity in the fiber of “aerospace anti-corrosion agent” ($K_{\text{source}}$). The discovery of its utility in household applications was a recognition of its high similarity to the nascent concept of a “general-purpose lubricant” ($K_{\text{target}}$), making it the first entity to populate that conceptual fiber and create a new market. At an abstract level, creativity can also be characterized as re-contextualization: an entity in the fiber of a source concept is recognized as crossing the threshold for a target concept, thereby making the new fiber non-empty for the first time.
6 Interpretation in Machine Learning
Similarity Field Theory provides a new geometric lens for understanding typical machine learning models. As discussed earlier in our “Beyond a purely statistical view” subsection, our aim here is not to rehash frequency co-occurrence arguments but to treat networks as knowledge structures of concept fibers; we instantiate this perspective with two canonical cases: neural networks and LLMs [5].
Neural networks as composed similarity fields
Consider a neural network trained to classify images of a cat. Let $K_{\text{Cat}}$ be the concept “cat”. The trained network can be viewed as approximating a similarity field
$S_{\text{NN}}(E_{\text{input}},K_{\text{Cat}})\in[0,1],$
where $E_{\text{input}}$ is an input entity such as a vector of pixel values.
Internal activations are typically unbounded real values. For interpretation, we fix a strictly increasing function $\varphi:\mathbb{R}\to[0,1]$ and define a calibrated activation $a^{\sim}=\varphi(a)\in[0,1]$. This allows us to read a neuron’s response as a similarity value without altering the network’s forward pass. A single neuron in the first hidden layer thus induces a primitive field
$S_{\text{neuron}}(E_{\text{input}},K_{\text{feature},w}):=\varphi(a(E_{\text{input}})),$
where $K_{\text{feature},w}$ denotes the latent feature concept associated with the neuron’s weight vector $w$. For self-similarity we explicitly normalise $S_{\text{neuron}}(K_{\text{feature},w},K_{\text{feature},w})=1$. In practice one can pick a prototype input $E^{\star}$ whose calibrated activation $\varphi(a(E^{\star}))$ is close to the supremum and interpret it as approximately achieving similarity $1$.
At a suitable level of abstraction, the output of a layer (or block) can be viewed as applying a map $g_{\ell}:[0,1]^{m}\to[0,1]$ to a vector of calibrated inputs, where $g_{\ell}$ is coordinatewise monotone and satisfies $g_{\ell}(1,\dots,1)=1$ on chosen prototypes. We do not require each primitive operation inside the layer to be a similarity field by itself; rather, after calibration we interpret the layer-level map as an aggregator over upstream field values. In common architectures this aggregator is implemented by nonnegative weighted sums, convex combinations, or pointwise products of calibrated activations, which keep the range in $[0,1]$ and preserve the prototype value $1$. Iterating this construction over layers yields, at the level of layer outputs, a composed similarity field whose final readout approximates $S_{\text{NN}}(E_{\text{input}},K_{\text{Cat}})$.
Training can then be viewed as similarity maximization: the network parameters are updated so that entities belonging to the cat fiber $F_{\alpha}(K_{\text{Cat}})$ receive higher values of $S_{\text{NN}}(E_{\text{input}},K_{\text{Cat}})$, subject to task constraints. In this view, backpropagation shapes an internal similarity field rather than just fitting an input–output statistic.
$E_{\text{input}}$$K_{\text{feature},\mathbf{w}}$$S(E_{\text{input}},K_{\text{feature},\mathbf{w}})$
Figure 2: A single neuron as a primitive similarity field.
$E_{\text{input}}$$K_{L1,1}$$K_{L1,2}$$K_{L1,3}$$K_{L2,1}$$K_{L2,2}$$K_{L2,3}$$K_{\text{Cat}}$$S_{\text{NN}}(E_{\text{input}},K_{\text{Cat}})$
Figure 3: A neural network composes simple similarity fields into a complex one.
Language models as token-level similarity generators
At the language interface, each vocabulary token can be treated as a concept $K_{\text{token}}\in U$, and an input prompt as a finite sequence of such token-concepts whose composition induces a higher-order concept $K_{\text{prompt}}$. The trained language model itself is an entity $K_{\text{Language}}$ that represents a vast meta-concept of language and, simultaneously, a generative operator $G$. For a given token concept $K_{\text{token}}$, its fiber $F_{\alpha}(K_{\text{token}})$ collects entities in $U$ that align with the token’s usage in human language and culture at threshold $\alpha$. When $G$ acts on a prompt, the prediction head compares the synthesized prompt concept $K_{\text{prompt}}$ against every token concept in its vocabulary; the resulting normalized probability distribution over tokens can be interpreted, after calibration, as a monotone surrogate of the similarity values $S(E,K_{\text{prompt}})$. Selecting a high-probability next token $E^{\prime}_{\text{next}}$ therefore approximately enforces $S(E^{\prime}_{\text{next}},K_{\text{prompt}})\geq\alpha$ for some effective threshold $\alpha$, so that the model behaves as a generative operator that tends to remain within the fiber of the prompt concept under the Similarity Field Theory view.
A path towards interpretability
In the SFT view, a trained network is a composition of calibrated similarity fields and can be analysed in terms of the concepts attached to its units. Each neuron computes a value $S_{\text{neuron}}(E_{\text{input}},K_{\text{latent}})$, which we interpret as the similarity between an input entity and a latent concept $K_{\text{latent}}$. In an idealised limit, the concept represented by that neuron is fully captured by its core fiber: the set of all inputs that drive its calibrated activation to the maximum value $1$,
$F_{1}(K_{\text{latent}})=\{\,E_{\text{input}}\in U\mid S_{\text{neuron}}(E_{\text{input}},K_{\text{latent}})=1\,\}.$
Recovering, even approximately, the inputs in this fiber amounts to recovering what the neuron “means”. From this perspective, interpretability becomes a geometric problem: decomposing the network into its constituent conceptual fibers and understanding how these fibers are composed across layers.
As Similarity Inference
This interpretation suggests a deeper principle. Entities we perceived, in this view, are not fundamental givens; they are stable structures inferred from a dynamic field of perceived similarities. We contend that we never perceive an entity like a “cat” directly; what we perceive is a coherent pattern of similarities in the spectrum of light intensities reaching our sensory organs. Even the perception of a fundamental quantity like length is an act of relational inference. When observing a 100-meter building, we do not directly apprehend the physical quantity of “100 meters.” This makes a framework centered on similarity not just a useful model for machine learning, but a reflection of the very nature of cognition itself.
Relations to other formalisms
Similarity as relational inference
SFT frames perception and judgement primarily as relational inference on a dynamic similarity field, without presupposing direct access to objects. In this view, stable entities appear as persistent superlevel-set structures, i.e., regularities in similarity relations that remain coherent under small perturbations. This perspective complements rather than replaces raw frequency-based descriptions by providing a geometric language for the structure that underwrites stable readouts.
Analogical reasoning
Analogical judgments can be expressed as structure-preserving maps between fibers. Given concepts $K_{\text{src}}$ and $K_{\text{tgt}}$, an analogy aligns subsets where $S(\cdot,K_{\text{src}})$ and $S(\cdot,K_{\text{tgt}})$ share level-set structure. This fiber-consistent view turns analogical correspondences into constraints on similarity geometry and permits quantitative tests by checking whether mapped instances remain within target fibers.
Symbolic constraints and fuzzy decisions
Symbolic constraints can be cast as predicates on fibers. A discrete constraint corresponds to membership in an intersection of fibers at prescribed thresholds, while graded decisions correspond to $\alpha$-level choices on a single fiber. This recovers classical fuzzy $\alpha$-cuts within SFT, with similarity $S(\cdot,K)$ serving as the membership function and $F_{\alpha}(K)$ as its superlevel set. In this way, crisp logic, soft decisions, and their mixtures can be expressed within the same similarity-field framework, linking rule-like reasoning with graded conceptual membership [18].
Neural components under calibrated similarity
With calibrated readouts, neural networks can be viewed as hierarchical compositions of primitive similarity fields, where intra- and inter-layer aggregations act as approximately monotone operators. Under this lens, mechanistic analyses that recover circuits, features, or probes correspond to identifying latent concepts $K$ and characterizing their fibers. Stable predictions arise when the determining vectors anchor or become confined to level sets of a readout, consistent with the Stability Theorem. This perspective places structural interpretability and readout stability within a single geometric language.
7 Conclusion
Similarity Field Theory treats dynamic similarity as the primitive for cognition and intelligence. By relaxing the strict axioms of classical mathematics and formalizing entities, fibers (as superlevel sets), and generative operators, we provide a precise language to describe the evolution of conceptual structures via sequences. Within this language, we proved two foundational principles: the Incompatibility Theorem, which reveals the necessary consequences of asymmetry, and the Stability Theorem, which establishes the causal dependency of stable outcomes on stable foundations through anchors or eventual confinement to a level set. The framework culminates in a generative definition of intelligence as the preservation of these structures—an operational principle that is both generalizable and measurable. Networks can be analyzed as compositions of similarity fields, enabling a principled route to interpretability. In short, this framework shifts the center of gravity from statistics to the geometry of similarity: intelligence as generative fiber preservation, and stability as eventual confinement within level sets. Ultimately, this work points to a shift in perspective—from objects to similarity relations—offering a computable language for modeling cognitive structure.